OpenStudy (anonymous):

Help! Write the equation of a line through C that it parallel to AB Writhe the equation of a line through B that is perpendicular to AC PIC BELOW

OpenStudy (anonymous):



OpenStudy (anonymous):

@shubhamsrg think you could help me out...sorry if I tag u in alota of questions

OpenStudy (shubhamsrg):

what is the slope of line AB ?

OpenStudy (anonymous):

slope is x2-x1 y2-y1 right I cant remember the formula right now

OpenStudy (shubhamsrg):

slope is ( y2 -y1 )/(x2 -x1) plug in values, and tell me, what is the slope ?

OpenStudy (anonymous):

AB=0.17

OpenStudy (anonymous):

I meant 0.75

OpenStudy (shubhamsrg):

yes, 0.75 is right now, equation of a line passing through given point and known slope is (y-y1) = m(x-x1) plug in m=0.75 x1 = 5 y1 = -1

OpenStudy (anonymous):

im a little lost now... -1-5=0.74-5

OpenStudy (shubhamsrg):

hmm, no it will be (y- (-1) ) = 0.75 (x - 5)

OpenStudy (anonymous):

that's the equation??

OpenStudy (anonymous):

or answer lol

OpenStudy (shubhamsrg):

yes.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

what about the equation of a line through b that is perpendicular to AC do the same thing only use the slope for AC?

OpenStudy (shubhamsrg):

yes, use the slope of AC. this time it is perpendicular, so you'll have to multiply the slope by -1and take its reciprocal. might sound bit complicated but is easy

OpenStudy (anonymous):

ok so slope of AC is 5 so it would be (y-(-1)=5(x-5)?

OpenStudy (shubhamsrg):

slope of AC is not 5, recheck

OpenStudy (anonymous):

is it -7

OpenStudy (shubhamsrg):

close, but not exactly right. give it another try :)

OpenStudy (anonymous):

7/-1

OpenStudy (shubhamsrg):

nopes, it should be -1/7

OpenStudy (anonymous):

ooh ok I got it I did it upside down

OpenStudy (anonymous):

but then I do the same thing like I put above only with a 7

OpenStudy (shubhamsrg):

nvm, so now this is perpendicular, so you revrse this and take its -ve reverse = -7 negative = 7

OpenStudy (anonymous):

ok I get theat

OpenStudy (shubhamsrg):

now, this line passes through (2,3) so eqn will be (y-3) = 7(x-2)

OpenStudy (anonymous):

oooh ok